# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ Abstract
Policy-relevant treatment effect estimation in a marketplace setting requires taking into account both the direct benefit of the treatment and any spillovers induced by changes to the market equilibrium. The standard way to address these challenges is to evaluate interventions via cluster-randomized experiments, where each cluster corresponds to an isolated market. This approach, however, cannot be used when we only have access to a single market (or a small number of markets). Here, we show how to identify and estimate policy-relevant treatment effects using a unit-level randomized trial run within a single large market. A standard Bernoulli-randomized trial allows consistent estimation of direct effects, and of treatment heterogeneity measures that can be used for welfare-improving targeting. Estimating spillovers—as well as providing confidence intervals for the direct effect—requires estimates of price elasticities, which we provide using an augmented experimental design. Our results rely on all spillovers being mediated via the (observed) prices of a finite number of traded goods, and the market power of any single unit decaying as the market gets large. We illustrate our results using a simulation calibrated to a conditional cash transfer experiment in the Philippines.
@@ 1 Introduction
In this paper, we introduce a framework for treatment effect estimation in a market equilibrium setting without relying on parametric, distributional, or homogeneity assumptions. Rather than relying on market splitting, we leverage a structural assumption whereby all spillovers are mediated through the (observed) equilibrium prices of a finite number of traded goods. Our main finding is that integrating such price equilibrium effects into the potential-outcomes model for causal inference enables us to estimate policy-relevant treatment effects using unit-level experiments. In particular, we show how our approach yields asymptotic characterizations of policy-relevant average and conditional average treatment effects, and motivates the design of new forms of randomized trials and targeting rules.
@@
We frame our model in terms of a potential outcomes specification that allows for cross-unit interference (Hudgens & Halloran 2008, Manski 2013, Aronow & Samii 2017). When equilibrium effects are removed, the model reduces to the standard Neyman-Rubin potential outcomes framework (Imbens & Rubin 2015). In a market with $n$ participants, an intervention is assigned randomly conditional on pre-treatment covariates. Each individual’s production and consumption choices are determined by latent supply and demand functions that may be shifted by the intervention (Angrist et al. 2000, Heckman & Vytlacil 2005). Units interact via a market modeled as a multiple goods economy in general equilibrium, where individuals’ production and consumption decisions depend only on their own type, their own treatment, and their expectation of the market price (Walras 1900, Mas-Colell et al. 1995). In this setup, equilibrium prices induce an exposure mapping in the sense of Manski (2013) and Aronow & Samii (2017): One unit’s treatment impacts another’s outcomes only through the treatment’s impact on the price equilibrium, which rules out peer effects or other forms of network-type interference. The price ensures that average net demand (demand minus supply) is equal to zero. The validity of this restriction on spillovers is context-specific, and depends both on the nature of the intervention and the type of market.
@@
We start our analysis by considering the marginal effect of increasing treatment probabilities for all market participants; following Carneiro et al. (2010), we refer to this quantity as a marginal policy effect. As shown in Hu et al. (2022), the marginal policy effect can be disposed into a direct and indirect treatment effect, where the indirect effect is a general measure of spillover effects. We then analyze the sample average direct and indirect effects of a binary treatment in equilibrium (Halloran & Struchiner 1995, Hu et al. 2022, Sävje et al. 2021). In finite samples, these estimands are difficult to characterize statistically because the price equilibrium induces a complex form of data dependence. We show, however, that under an assumption that the $n$ participants in the market are independently sampled from a population distribution, we can use techniques based on empirical-process theory to derive simple large-sample limiting expressions for the direct and indirect effects as $n$ grows. These limiting expressions can then be used to guide estimation and inference for both sample and population treatment effects.
@@ 1.1 Related Work
Our model combines a potential outcomes model with some limited structural assumptions on the price equilibrium. The insight that structural modeling can be used to address endogeneity issues in causal inference has a long tradition in economics (Heckman 1979, Roy 1951, Heckman et al. 1998), and there is a large literature that uses structural modeling to develop econometric methods for causal inference in models with endogenous treatment selection (Angrist et al. 2000, Carneiro et al. 2011, Heckman & Vytlacil 2005). To date, however, there has been less emphasis on methods for estimating equilibrium effects in systems with randomized trials, where treatment selection is not an issue.
We do note a growing literature that combines data from a randomized trials with parametric structural models to estimate both partial and general equilibrium policy effects. For example, Neilson et al. (2019) uses a randomized control trial to estimate the direct effect of information provision on school choice but simulates a fully parametric structural model to evaluate equilibrium effects of the treatment. In contrast, our approach uses a more complex unit-level experiment and the structure of the price equilibrium to analyze equilibrium effects without imposing a parametric model. To accomplish this, we rely on a representation of population estimands as a combination of price elasticities and expectations of the outcome and demand distribution. This is related to the sufficient statistics approach, as described in Chetty (2009), which also expresses policy effects in terms of combinations of estimable elasticities without imposing a specific parametric model, although this is usually in a partial equilibrium setting. In the macroeconomics literature, Wolf (2023) uses a semi-structural approach to decompose the effect of economic shocks into partial equilibrium effects and general equilibrium effects (e.g., price effects), but uses time series methods to estimate the price effects, in contrast to the unit-level approach presented in this paper.
@@ 2 Model and Estimands
We work under a potential-outcomes model of a market with $n$ participants, where each participant is drawn independently from a population. A binary intervention affects individual supply, demand and outcomes in the market. We use this model to define the estimands of interest for our paper, both at the sample level and at the population level. We use “sample” to refer to a market with $J$ goods and $n$ participants, since we assume that the observed equilibrium arises in this market and in many cases we will observe outcomes for all $n$ market participants. However, as we explain in more detail in Section 3, our inference approach will also allow for settings where we can only observe outcomes for a sub-sample of the $n$-sized market (but equilibrium effects are still determined by the full size-$n$ sample).
In a market of size $n$, each market participant $i=1,\,\ldots,\,n$ is characterized by an outcome function $Y_{i}(w_{i},\,p)$, which is a map from $\{0,1\}\times\mathbb{R}^{J}\to\mathbb{R}$, a net demand function $Z_{i}(w_{i},\ p)$, which is a map from $\{0,1\}\times\mathbb{R}^{J}\to\mathbb{R}^{J}$, and covariates $X_{i}\in\mathcal{X}$. These quantities are drawn jointly from a population distribution. Net demand is defined as demand minus supply. We also allow for an exogenous individual-level price shifter $U_{i}\in\mathbb{R}^{J}$. For our analysis of classical randomized trials, $U_{i}=0$, but we will also consider a class of experiments that randomize discounts or fees in a market, where $U_{i}\neq 0$ and is also exogenous. Each market participant is also assigned a treatment $W_{i}\in\{0,1\}$. Using these variables, we can now define potential outcomes in our model. Throughout this paper, we use bold-face notation to denote a vector-valued quantity collected across all market participants; for example, we write $\mathbf{Z}=(Z_{1},\ \ldots,\,Z_{n})$.
@@
is non-empty with probability at least $1-e^{-cn}$ for all $n$. On the event where this set is non-empty, the market price is in this set, $P_{n}(\mathbf{w})\in\mathcal{S}_{\mathbf{w}}$.
In Section 3, we will use observations of net demand, outcomes, treatments and price-shifters in a single market of size $n$ to estimate policy-relevant treatment effects. Assumption 1 is the main structural assumption in the paper that makes this possible. Treatment and market prices directly impact outcomes and net demand for each individual. Prices are endogenous, and are determined by a sample equilibrium condition on net demand. Because of this endogeneity, there are spillovers from individual $i$ to individual $j$ through their effect on the equilibrium that the market reaches. Although our model allows for equilibrium spillovers, it rules out other types of spillovers, such as through peer effects or some other network mechanism. The validity of this restriction on interference primarily depends on the nature of the intervention and the context of the market. A subsidy allocated to suppliers in an online market may, by design, be impossible to share with those in the control group, so spillovers occur only through market prices. On the other hand, an information intervention that affects consumption in a village economy may have spillovers both through social networks in the village and through the market equilibrium.
The second part of the assumption characterizes how prices form in finite samples. The finite market price does not need to exactly clear the market; instead, we impose a restriction on the rate at which the error in the market-clearing condition shrinks with the sample size. Furthermore, it does not need to be the unique price meeting this condition, although we will require uniqueness at the population level, i.e., in the limit as the sample size grows large (Assumption 2). Proposition 21 in Appendix D provides explicit sufficient conditions for Assumption 1 to hold in a single good market without price shifters and with discontinuous net demand functions.
For a market with $n$ participants where a vector of treatments $\bm{W}$ and matrix of price perturbations $\mathbf{U}$ is realized, we observe $Y_{i}=Y_{i}(\bm{W})$ and we also observe realized net demand $Z_{i}=Z_{i}(W_{i},\bm{P}_{n}(\bm{W})+U_{i})$. For the majority of the paper, we will consider settings without price shifters (i.e., with $\mathbf{U}=0$), where treatments are Bernoulli-randomized in the sample as described below. In Section 3, we will also consider a class of augmented randomized trials with exogenous price shifters $U_{i}$ whose size shrinks as the market size grows large (Design 2).
@@
Bernoulli Randomized Trial. Each unit has a known randomization probability $\pi_{i}$ satisfying $\eta\leq\pi_{i}\leq 1-\eta$ for some $\eta>0$, such that treatment is generated as $W_{i}\sim\text{Bernoulli}(\pi_{i})$. The $\pi_{i}$ may be constant, functions of $X_{i}$, or randomly generated. There are no price perturbations, so $U_{i}=0$ for all $i\in\{1,\,\ldots,n\}$.
@@
$\tau_{\text{MPE}}=\frac{d}{d\varepsilon}V(\bm{\pi}+\bm{\varepsilon})|_{\varepsilon=0}.$ (4)
@@
The estimands $\tau_{\text{ADE}}$ and $\tau_{\text{AIE}}$ are defined at the sample level. We also find it useful to introduce population versions of our target estimands, and define the following whenever the limits below exist:
@@
In the next section, we will show that under our model and an additional set of regularity conditions, these limits in fact exist and the population estimands have a tractable representations. These representations will then be useful for designing estimators that are consistent for both sample and population-level targets. In addition, just as in the literature on causal inference under SUTVA, in some settings the population estimand is the more relevant target. For example, we may have data from a small market and are interested in doing inference on a similar but larger market, in a setting where there is significant homogeneity across markets.
@@ 2.1 Large Sample Characterization
We start by making regularity assumptions on the expected net demand, and in particular assume that there is a unique equilibrium price in the population. Market values will be in a compact set as long as all net demand functions are weakly negative for prices above an upper bound and weakly positive for prices smaller than some lower bound. In a single good market, uniqueness of $p^{*}_{\pi}$ only requires continuity and strict monotonicity of $z(p)$ in $p$. In a market with multiple goods, uniqueness and existence of the equilibrium price can be shown by ensuring the expected net demand function is a contraction, as is common in the literature on models of strategic behavior (Cornes et al. 1999, Van Long & Soubeyran 2000). Other approaches that rely on more primitive assumptions are also possible; under a zero-degree homogeneity assumption in net demand then a gross-substitutes condition ensures uniqueness (Arrow & Hahn 1971).
@@
Market prices take values in a compact set $\mathcal{S}\subset\mathbb{R}^{J}$ almost surely. Given any randomization policy $\pi$, there is a unique population market-clearing price $p^{*}_{\pi}\in\mathcal{S}$ that satisfies $z(p^{*}_{\pi})=0.$ Let the Jacobian $\xi_{z}=\nabla_{p}z(p^{*}_{\pi})$, where row $j$ of the $J\times J$ matrix is $\nabla^{\top}_{p}z_{j}(p^{*}_{\pi})$. The Jacobian is full rank.
@@
Assumption 4 allows the outcome function to be the sum of a term $H_{i}(w,\,p)$ with a Lipschitz dependence on prices, and a bilinear function of prices and net demand. Here, $H_{i}(w,\,p)$ could be a direct measurement of welfare, which under standard rational choice assumptions should vary continuously with prices. Or, it could be any Lipschitz transformation of net demand and prices when net demand is itself Lipschitz. When net demand is not smooth at the individual level, we allow for a bilinear function of $p$ and $Z_{i}(w,\,p)$, which captures natural researcher-constructed outcomes such as surplus or profit. This allows for outcomes that are discontinuous in prices, but the discontinuity must only through simple transformations of net demand. We note that all our results below also hold if $Y_{i}(w,\,p)$ depends on $\left(\alpha+\beta^{\top}p\right)G_{i}(w,\,p)$ for any function $G_{i}(w,\,p)$ satisfying Assumption 3; however, for simplicity, we do not add this degree of flexibility in the notation. [^2] This captures researcher-constructed outcomes that depend on possibly discontinuous consumption or production of certain goods in a monotonic way. For example, in Section 5, health outcomes depend on consumption of protein sources, and this relationship could be non-linear in our framework, even when individual-level consumption is discontinuous.
@@
The outcome functions can be decomposed as the sum of a Lipschitz welfare function and a (random) function of $p$ times net demand,
$Y_{i}(w,\,p)=H_{i}(w,\,p)+\left(\alpha+\beta^{\top}p\right)\cdot Z_{i}(w,\,p),$ (8)
where $H_{i}(w,\,p)$ is $M$-Lipschitz in $p$, $\alpha\in\mathbb{R}^{J}$ and $\beta\in\mathbb{R}^{J\times J}$.
Finally, in Assumption 5 we list some additional regularity assumptions, which are standard in the literature on asymptotic statistics. Although individual net demand and outcome functions may be discontinuous, we require their expectation to vary smoothly in prices. The weak continuity assumption effectively requires that discontinuity points cannot concentrate at specific values of $p$; for example, in a one-dimensional threshold-crossing model, our weak continuity assumption is equivalent to the cutoff point distribution having a density that is uniformly bounded from above.
@@
Our first result is that, given these assumptions, the sample equilibrium price concentrates on $p^{*}_{\pi}$ as the sample size grows, and has an asymptotically linear representation (i.e., to first order, random price fluctuations can be written as a sum of additive contributions from each unit). All proofs are given in the appendix. Our technical results rely heavily on asymptotic equicontinuity results from empirical process theory as described in van der Vaart & Wellner (1996).
@@
Under Assumptions 1, 2, 3 and 5, the finite sample equilibrium price under Design 1 and Design 2 satisfies
$P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-1/2}),$ (9)
and furthermore $\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\xi_{z}^{-1}\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\left(\xi_{z}^{-1}\right)^{\top}\right)$, where $\xi_{z}$ is as defined in Assumption 2.
This result is a crucial building block towards the rest of the theory in the paper, since it provides a representation of $P_{n}(\bm{W})$ in terms of the fixed price $p^{*}_{\pi}$ and an average of IID terms. This will be helpful in characterizing the asymptotic variance of various random variables in our model. Before introducing our estimators in Section 3, we next apply this result to characterizing the relationship between the sample and population-level average direct and indirect effect. To this end, we first show that $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ exist and have a simple representation in terms of moments of the population distribution.
@@
Under Assumptions 1, 2, 3, 4 and 5, the population-level estimands in the Bernoulli Design 1 are:
@@
$\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\Big{(}Y_{i}(1,p^{*}_{\pi})-Y_{i}(0,p^{*}_{\pi})-\left(\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})+(1-\pi_{i})\Delta_{i}(1,\,p^{*}_{\pi})\right)\Big{)}+o_{p}(1/\sqrt{n}),$ (11)
@@
Meanwhile, for the indirect effect, $\tau_{\text{AIE}}$ converges to $\tau_{\text{AIE}}^{*}$, but Theorem 4 does not provide a rate of convergence. We conjecture that establishing useful rates of convergence here would require replacing Assumptions 2 and 3 with stronger assumptions that are more explicit about finite-sample price formation. Under our current assumptions, the convergence rate of $\tau_{\text{AIE}}$ can be slower than $\sqrt{n}$; see Appendix D.2 for additional explanation in the context of a simple example. Since our estimator in Section 3 targeting $\tau_{\text{AIE}}^{*}$ also converges at a slower than $\sqrt{n}$ rate, we conjecture that confidence intervals built for $\tau^{*}_{\text{AIE}}$ will also generally cover $\tau_{\text{AIE}}$; we also note that this holds in our numerical experiments.

### v4

@@ Abstract
Policy-relevant treatment effect estimation in a marketplace setting requires taking into account both the direct benefit of the treatment and any spillovers induced by changes to the market equilibrium. The standard way to address these challenges is to evaluate interventions via cluster-randomized experiments, where each cluster corresponds to an isolated market. This approach, however, cannot be used when we only have access to a single market (or a small number of markets). Here, we show how to identify and estimate treatment effects using a unit-level randomized trial run within a single large market. A standard Bernoulli-randomized trial allows consistent estimation of direct effects, and of treatment heterogeneity measures that can be used for welfare-improving targeting. Estimating spillovers—as well as providing confidence intervals for the direct effect—requires estimates of price elasticities, which we provide using an augmented experimental design. Our results rely on all spillovers being mediated via the (observed) prices of a finite number of traded goods, and the market power of any single unit decaying as the market gets large. We illustrate our results using a simulation calibrated to a conditional cash transfer experiment in the Philippines.
@@ 1 Introduction
In this paper, we introduce a framework for treatment effect estimation in a market equilibrium setting without relying on parametric, distributional, or homogeneity assumptions. Rather than relying on market splitting, we leverage a structural assumption whereby all spillovers are mediated through the (observed) equilibrium prices of a finite number of traded goods. Our main finding is that integrating such price equilibrium effects into the potential-outcomes model for causal inference enables us to estimate spillover-aware treatment effects using unit-level experiments. In particular, we show how our approach yields asymptotic characterizations of policy-relevant average and conditional average treatment effects, and motivates the design of new forms of randomized trials and targeting rules.
@@
We frame our model in terms of a potential outcomes specification that allows for cross-unit interference (Hudgens & Halloran 2008, Manski 2013, Aronow & Samii 2017). When equilibrium effects are removed, the model reduces to the standard Neyman-Rubin potential outcomes framework (Imbens & Rubin 2015). In a market with $n$ participants, an intervention is assigned randomly conditional on pre-treatment covariates. Each individual’s production and consumption choices are determined by latent supply and demand functions that may be shifted by the intervention (Angrist et al. 2000, Heckman & Vytlacil 2005). Units interact via a market modeled as a multiple goods economy in general equilibrium, where individuals’ production and consumption decisions depend only on their own type, their own treatment, and their expectation of the market price (Walras 1900, Mas-Colell et al. 1995). In this setup, equilibrium prices induce an exposure mapping in the sense of Manski (2013) and Aronow & Samii (2017): One unit’s treatment impacts another’s outcomes only through the treatment’s impact on the equilibrium price, which rules out peer effects or other forms of network-type interference. The price ensures that average net demand (demand minus supply) is equal to zero. The validity of this restriction on spillovers is context-specific, and depends both on the nature of the intervention and the type of market.
@@
Our analysis begins by considering the marginal effect of increasing treatment probabilities for all market participants; following Carneiro et al. (2010), we refer to this quantity as a marginal policy effect. As shown in Hu et al. (2022), the marginal policy effect can be decomposed into a direct and indirect treatment effect, where the indirect effect is a general measure of spillover effects. We then analyze the sample average direct and indirect effects of a binary treatment in equilibrium (Halloran & Struchiner 1995, Hu et al. 2022, Sävje et al. 2021). In finite samples, these estimands are difficult to characterize statistically because the price equilibrium induces a complex form of data dependence. We show, however, that under an assumption that the $n$ participants in the market are independently sampled from a population distribution, we can use techniques based on empirical-process theory to derive simple large-sample limiting expressions for the direct and indirect effects. These limiting expressions can then be used to guide estimation and inference for both sample and population treatment effects.
@@ 1.1 Related Work
Our model combines a potential-outcomes model with some limited structural assumptions on the price equilibrium. The insight that structural modeling can be used to address endogeneity issues in causal inference has a long tradition in economics (Heckman 1979, Heckman et al. 1998, Roy 1951), and there is a large literature that uses structural modeling to develop econometric methods for causal inference in models with endogenous treatment selection (Angrist et al. 2000, Carneiro et al. 2011, Heckman & Vytlacil 2005). To date, however, there has been less emphasis on methods for estimating equilibrium effects in systems with randomized trials, where treatment selection is not an issue.
We do note a growing literature that combines data from a randomized trials with parametric structural models to estimate both partial and general equilibrium policy effects. For example, Allende et al. (2019) uses a randomized control trial to estimate the direct effect of information provision on school choice but simulates a fully parametric structural model to evaluate equilibrium effects of the treatment. In contrast, our approach uses a more complex unit-level experiment and the structure of the price equilibrium to analyze equilibrium effects without imposing a parametric model. To accomplish this, we rely on a representation of population estimands as a combination of price elasticities and expectations of the outcome and demand distribution. This is related to the sufficient statistics approach, as described in Chetty (2009), which also expresses policy effects in terms of combinations of estimable elasticities without imposing a specific parametric model, although this is usually in a partial equilibrium setting. In the macroeconomics literature, Wolf (2023) uses a semi-structural approach to decompose the effect of economic shocks into partial equilibrium effects and general equilibrium effects (e.g., price effects), but uses time series methods to estimate the price effects, in contrast to the unit-level approach presented in this paper.
@@ 2 Model and Estimands
We first introduce a potential-outcomes model of a market with $n$ participants, where each participant is drawn independently from a population. A binary intervention affects individual supply, demand and outcomes in the market. We use this model to define the estimands of interest for our paper, both at the sample level and at the population level. We use “sample” to refer to a finite-sized market with $J$ goods and $n$ participants. Throughout this section, we will refer to two example interventions: The first is a supplier-level subsidy in an online market setting; the second is a household-level cash transfer that affects economic choices in a village economy.
[…]
Assumption 1.
[…]
@@
is non-empty with probability at least $1-e^{-c_{1}n}$ for all $n$. On the event where this set is non-empty, the market price is in this set, $P_{n}(\mathbf{w})\in\mathcal{S}_{\mathbf{w}}$.
In Section 3, we will use observations of net demand, outcomes, treatments and price-shifters in a single market of size $n$ to estimate policy-relevant treatment effects. Assumption 2 is the main structural assumption in the paper that makes this possible. Treatment and market prices directly impact outcomes and net demand for each individual. Prices are endogenous, and are determined by a sample equilibrium condition on net demand. Because of this endogeneity, there are spillovers from individual $i$ to individual $j$ through their effect on the equilibrium that the market reaches. Although our model allows for equilibrium spillovers, it rules out other types of spillovers, such as through peer effects or some other network mechanism. The validity of this restriction on interference primarily depends on the nature of the intervention and the context of the market. A subsidy allocated to suppliers in an online market may, by design, be impossible to share with those in the control group, so spillovers occur only through market prices. On the other hand, an intervention that affects consumption in a village economy may have spillovers both through social networks in the village and through the market equilibrium.
The second part of the assumption characterizes how prices form in finite samples. The finite market price does not need to exactly clear the market; instead, we impose a restriction on the rate at which the error in the market-clearing condition shrinks with the sample size. Furthermore, it does not need to be the unique price meeting this condition, although we will require uniqueness at the population level, i.e., in the limit as the sample size grows large (Assumption 3). [^4] Proposition 23 in Appendix D provides explicit sufficient conditions for Assumption 2 to hold in a single good market without price shifters and with discontinuous net demand functions.
We will start by considering results that hold under a standard Bernoulli-randomized RCT (Design 1). In Section 3, we will show how augmenting Bernoulli-randomized trial with random individual-level price shifters (Design 2) enables further estimation and inferential results. Design 2 will discussed in more detail in Section 3.2.
@@
Bernoulli Randomized Trial. Each unit has a known randomization probability $\pi_{i}$ satisfying $\eta\leq\pi_{i}\leq 1-\eta$ for some $\eta>0$, and treatment is generated as $W_{i}\sim\text{Bernoulli}(\pi_{i})$. The $\pi_{i}$ may be constant, functions of $X_{i}$, or random. [^5] In Design 1 there are no price perturbations, i.e., $U_{i}=0$ for all $i\in\{1,\ldots,n\}$.
Design 2.
Augmented Randomized Trial. Treatment $W_{i}$ is randomized according to Design 1. In addition, the experimenter generates random price perturbations $U_{i}\in\mathbb{R}^{J}$, such that for $i\in\{1,\,\ldots,n\}$ and $j\in\{1,\,\ldots,J\}$, $U_{ij}$ is drawn independently and uniformly at random from $\{-h_{n},+h_{n}\}$ where $h_{n}=cn^{-\alpha}$ with $\frac{1}{4}<\alpha<\frac{1}{2}$ and $c>0$ is a constant.
@@
$\tau_{\text{MPE}}=\frac{d}{d\varepsilon}V(\bm{\pi}+\varepsilon\mathbf{1}_{n})|_{\varepsilon=0}.$ (4)
@@
The estimands $\tau_{\text{ADE}}$ and $\tau_{\text{AIE}}$ are defined at the sample level. We also find it useful to introduce population versions of our target estimands, and define the following whenever the limits below exist. $\mathbb{E}[\cdot]$ reflects expectations over all sources of randomness in our model: the sampling from the population distribution in Assumption 1, the price-shifter assignment process, and the treatment assignment process.
@@
In the next section, we will show that under our model and an additional set of regularity conditions, these limits in fact exist and the population estimands have a tractable representations. These representations will then be useful for designing estimators that are consistent for both sample and population-level targets. In addition, just as in the literature on causal inference under SUTVA, in some settings the population estimand is the more relevant target. For example, we may have data from a small market and are interested in doing inference on a similar but larger market.
@@ 2.1 Large Sample Characterization
We start by making regularity assumptions on the expected net demand, and in particular assume that there is a unique equilibrium price in the population. Market prices will be in a compact set as long as all net demand functions are weakly negative for prices above an upper bound and weakly positive for prices smaller than some lower bound. In a single good market, uniqueness of $p^{*}_{\pi}$ only requires continuity and strict monotonicity of $z(p)$ in $p$. In a market with multiple goods, uniqueness and existence of the equilibrium price can be shown by ensuring the expected net demand function is a contraction, as is common in the literature on models of strategic behavior (Cornes et al. 1999, Van Long & Soubeyran 2000). Other approaches that rely on more primitive assumptions are also possible; for example, under a zero-degree homogeneity assumption in net demand a gross-substitutes condition ensures uniqueness (Arrow & Hahn 1971).
@@
Market prices take values in a compact set $\mathcal{S}\subset\mathbb{R}^{J}$ almost surely. Given any randomization policy $\pi$, there is a unique population market-clearing price $p^{*}_{\pi}\in\mathcal{S}$ that satisfies $z(p^{*}_{\pi})=0.$ The Jacobian $\xi_{z}=\nabla_{p}z(p^{*}_{\pi})$ (i.e., the $J\times J$ matrix with $j$-th row $\nabla^{\top}_{p}z_{j}(p^{*}_{\pi})$) is full rank.
@@
Under Assumption 5, the outcome function is the sum of an individual-specific random function that is Lipschitz in prices and an additional term that is Lipschitz in net demand. This allows for outcomes that are discontinuous in prices through net demand. Supplier profit, for example, is a special case of Assumption 5, where in the single-good case, profit is $Y_{i}(w,p)=-(p-\Gamma_{i}(w))Z_{i}(w,\,p)\mathbbm{1}(Z_{i}(w,\,p)<0)$ and $\Gamma_{i}(w)$ represents a firm’s production costs under treatment $w$. In Section 5, outcomes are a measurement of height for young children, and the treatment is a household-level cash transfer. Assumption 5 allows for a variety of flexible data-generating processes for height; one example is a random coefficients model of height that is polynomial in consumption, where $\Gamma_{i}(w)$ are the random coefficients and noise term in the model.
@@
The outcome function is the sum of a random Lipschitz function of $p$, $H_{i}(w,\,p)$, and a fixed transformation $\psi(\cdot)$ of $Z_{i}(w,\,p)$, prices $p$, and a (possibly unobserved) bounded random variable $\Gamma_{i}(w)\in\mathcal{G}\subset\mathbb{R}^{m}$, that is Lipschitz in each of its arguments:
$Y_{i}(w,\,p)=H_{i}(w,\,p)+\psi(\Gamma_{i}(w),\,Z_{i}(w,\,p),\,p).$ (8)
Finally, in Assumption 6 we list some additional regularity assumptions, which are standard in the literature on asymptotic statistics. Although individual net demand and outcome functions may be discontinuous, we require their expectation to vary smoothly in prices. The weak continuity assumption in Part 2 also limits individual-level discontinuity by requiring that discontinuity points cannot concentrate at specific values of $p$. In our profit example, these two smoothness assumptions require the distribution of costs at a firm-level to be sufficiently smooth, even if production is discontinuous in prices.
@@
Our first result is that, given these assumptions, the sample equilibrium price concentrates on $p^{*}_{\pi}$ as the sample size grows, and has an asymptotically linear representation (i.e., to first order, random price fluctuations can be written as a sum of additive contributions from each unit). We can write $P_{n}(\bm{W})$ as a method-of-moments estimator, so its asymptotic representation has the usual form of Z-estimators (van der Vaart & Wellner 1996). All proofs are given in the appendix.
@@
Under Assumptions 1, 2, 3, 4 and 6 and either Design 1 or 2, the equilibrium price satisfies
$\begin{split}&P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-1/2}),\\ &\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\xi_{z}^{-1}\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\left(\xi_{z}^{-1}\right)^{\top}\right),\end{split}$ (9)
where $\xi_{z}$ is as defined in Assumption 3.
This result is a crucial building block towards the rest of the theory in the paper, since it provides a representation of $P_{n}(\bm{W})$ in terms of the fixed price $p^{*}_{\pi}$ and an average of IID terms. This asymptotic representation is valid whether or not the Bernoulli-randomized trial is augmented with price perturbations. It will be helpful in characterizing the asymptotic variance of various random variables in our model. Before introducing our estimators in Section 3, we next apply this result to characterizing the relationship between the sample and population-level average direct and indirect effect. To this end, we first show that $\tau^{*}_{\text{ADE}}$ and $\tau^{*}_{\text{AIE}}$ exist and have a simple representation in terms of moments of the population distribution. Our technical results rely heavily on concentration results from empirical process theory as described in van der Vaart & Wellner (1996).
@@
Under Assumptions 1, 2, 3, 4, 5 and 6 and either Design 1 or 2, the population estimands are
@@
$\tau_{\text{ADE}}=\frac{1}{n}\sum\limits_{i=1}^{n}\Big{(}Y_{i}(1,p^{*}_{\pi})-Y_{i}(0,p^{*}_{\pi})-\left(\pi_{i}\Delta_{i}(1,\,p^{*}_{\pi})+(1-\pi_{i})\Delta_{i}(0,\,p^{*}_{\pi})\right)\Big{)}+o_{p}(1/\sqrt{n}),$ (11)
@@
Meanwhile, for the indirect effect, $\tau_{\text{AIE}}$ converges to $\tau_{\text{AIE}}^{*}$, but Theorem 4 does not provide a rate of convergence. We conjecture that establishing useful rates of convergence here would require replacing Assumptions 3 and 4 with stronger assumptions that are more explicit about finite-sample price formation. Under our current assumptions, the convergence rate of $\tau_{\text{AIE}}$ can be slower than $\sqrt{n}$; see Appendix D.3 for additional explanation in the context of a simple example. Since our estimator in Section 3 targeting $\tau_{\text{AIE}}^{*}$ also converges at a slower than $\sqrt{n}$ rate, we conjecture that confidence intervals built for $\tau^{*}_{\text{AIE}}$ will also generally cover $\tau_{\text{AIE}}$; we also note that this holds in our numerical experiments.